A solid cube whose edges are 12 cm each, has been cut by a plane such that the plane bisects three edges meeting at a corner. the surface area of the remaining part of cube after the small piece is removed is :

A)  81 + 10 root3

B) 810 + 10 root 3

C) 810 - 10 root 3

D) 810 + 18 root3

Grade:10

1 Answers

Ravi
9 years ago
Answer is D
Total SA before dissection.: 6a2= 864 ...(A)
Areaof the faces removed=3(1/2* 6*3)= 54 ...(B)
Area of the new open surface added after dissection to the bigger piece= Area of Equilateral triangle of side 6*21/2= 31/2*(6*21/2)2/4= 18*31/2 ...(C)
New total surface area= A-B+C